OpenStudy (anonymous):

Line AB contains (0, 4) and (1, 6) Line CD contains points (2, 10) and (-1, 4). Lines AB and CD are (4 points) Question 2 options: 1) parallel because the slopes are the same 2) perpendicular because the slopes are the same 3) parallel because the product of the slopes is -1 4) perpendicular because the product of the slopes is -1 *Help please

OpenStudy (anonymous):

they have the same slope

OpenStudy (anonymous):

Use the slope formula to find the slopes, which is y2-y1/x2-x1. First line has slope: 2/1 or 2. Second line has slope -6/-3, or 2. Since they have the same slope, they have to be parallel.

OpenStudy (anonymous):

A ? @M4thM1nd @ridoleor

OpenStudy (anonymous):

for Line AB we have: slope = (6-4)/(1-0) = 2 for line CB we have: slope = (10-4)/(2-(-1)) = 2 as they have the same slope the are parallel